OpenStudy (happyfeet8040):

Nick works two jobs to pay for college. He tutors for $15 per hour and also works as a bag boy for $8 per hour. Due to his class and study schedule, Nick is only able to work up to 20 hours per week but must earn at least $150 per week. If t represents the number of hours Nick tutors and b represents the number of hours he works as a bag boy, which system of inequalities represents this scenario?

OpenStudy (happyfeet8040):

t + b greater than or equal to 20 15t + 8b = 150 t + b less than or greater to 20 15t + 8b greater than or equal to 150 t + b less than or greater to 20 15t + 8b less than or greater to 150 None of the systems shown represent this scenario.

OpenStudy (happyfeet8040):

Someone please help! I will give a medal.

OpenStudy (imstuck):

Ok what you have to do here is relate the hours of tutoring plus the hours of bagging to the hours he is allowed to work, which is 20. Then you relate what he makes tutoring plus what he makes bagging to what he must earn which is 150. So the first equation relating the hours is this: \[t+b \le20\]What this says is that the hours of tutoring plus the hours of bagging has to be less than or equal to 20 hours.

OpenStudy (imstuck):

The other inequality relating the money earned and needed is this:\[15t+8b \ge 150\]This says that the 15 dollars per hour for tutoring plus the 8 dollars per hour for bagging must be at least 150 dollars. So that is your second inequality.

OpenStudy (imstuck):

So the solution should read (and I think you may have some typos in your choices): t + b is less than or equal to 20 and 15t+8b is greater than or equal to 150

OpenStudy (imstuck):

I think it's the second choice with a typo made on accident.

OpenStudy (imstuck):

If not, it is "none of the systems..." But I think it is supposed to be the second one.

OpenStudy (happyfeet8040):

I agree! It makes more sense now, thank you so much. :)

OpenStudy (happyfeet8040):

Can you help me with a couple more.

OpenStudy (happyfeet8040):

@IMStuck

OpenStudy (imstuck):

Sure. What do you have?

OpenStudy (happyfeet8040):

Sophia invested some money in a bank at a fixed rate of interest compounded annually. The equation below shows the value of her investment after x years: f(x) = 500(1.05)x What was the average rate of change of the value of Sophia's investment from the second year to the fourth year? 14.13 dollars per year 28.25 dollars per year 50.00 dollars per year 56.50 dollars per year

OpenStudy (imstuck):

Is that \[f(x)=500(1.05)^{x}\]or\[f(x)=500(1.05)x\]

OpenStudy (imstuck):

Important. They are very different from each other.

OpenStudy (happyfeet8040):

How are they different? Sorry, I'm trying to understand.. I'm not so great at math haha.

OpenStudy (imstuck):

it's the first one.

OpenStudy (imstuck):

Is it the first one? That is x as an exponent.

OpenStudy (imstuck):

I'm going with that one.

OpenStudy (imstuck):

The general formula is this:\[A=A _{0}(1+r)^{t}\]

OpenStudy (imstuck):

Going with that one and using the formula and x = 2, you have this\[A=500(1.05)^{2}\]and A = 551.25

OpenStudy (imstuck):

Now we will do it with 4 years:\[A=500(1.05)^{4}\]and A = 607.75

OpenStudy (imstuck):

Now subtract the first from the second, the lower from the higher, and the difference is 56.50 Do you see that answer in your choices?

OpenStudy (imstuck):

Are you there? I'm here to help if you need more...

OpenStudy (happyfeet8040):

Yes, I'm here. Sorry, I was doing something.

OpenStudy (happyfeet8040):

So, you think it's the first one? @IMStuck